# Aligning MLLM Benchmark With Human Preferences via Structural Equation Modeling

### v1

@@ 4.2 Main Results
In this section, we report the evaluation outcomes of various open-source and closed-source MLLMs on the Gold benchmark under the Main Experiment setting. Table 2 presents the detailed results, where models are assessed across three cognitive layers—Perception, Memory, and Reasoning—further subdivided into interpretable subtasks such as color recognition, OCR, basic medical knowledge, and domain-specific logical reasoning.
Overall Performance. GPT-4o, the only closed-source model in our comparison, achieves strong performance with an overall score of 165.7, particularly excelling in perception and memory tasks. However, top-tier open-source models such as InternVL3-78B (172.9) and Qwen2.5-VL-72B (157.6) actually surpass or closely approach GPT-4o in overall capability, suggesting that the performance gap between proprietary and open-source MLLMs is narrowing.
Capability-Level Comparison. Across all models, Perception tasks (Color, Count, OCR) exhibit the highest average scores, indicating that current MLLMs are well-optimized for low-level visual understanding. In contrast, Reasoning tasks—especially those requiring domain-specific knowledge (e.g., Electronics, Math, Economics) are the most challenging, with wide performance variability. This suggests that abstract and high-order reasoning remains a major bottleneck for multimodal intelligence.
Model Size and Performance. Larger model variants generally perform better, especially within the InternVL3 and Qwen2.5-VL families. For example, InternVL3 improves from 116.9 (2B) to 172.9 (78B), showing consistent gains across all layers. However, despite these improvements, performance on certain reasoning subfields (e.g., electronics, economics) still lags, indicating that increased model size alone does not fully bridge capability gaps.
4.3 SEM-based Benchmark Evaluation Result Analysis
As shown in Table LABEL:tab:diagnostic_results, Gold achieves the best balance among the three evaluation dimensions, with the highest average score (0.656) across all metrics. It combines broad cognitive coverage, strong task informativeness, and robust statistical validity, demonstrating the strength of theory-driven benchmark restructuring under the SEM-based framework. In contrast, some existing benchmarks show localized strengths but lack holistic consistency. Moreover, Gold shows the highest alignment with human ratings (Pearson $r=0.7359$), reflecting its superior interpretability and evaluation reliability.
@@ 4.3 SEM-based Benchmark Evaluation Result Analysis
Table 3: Diagnostic metrics for existing MLLM benchmarks across four dimensions.
Benchmark $D_{div}$ $TC$ $D_{valid}$ Avg
SEEDBenchV2 0.605 0.871 0.081 0.519
MVBench 0.094 0.584 0.522 0.400
MMEBench 0.549 0.861 0.420 0.610
MMMUBench 0.492 0.808 0.442 0.581
BLINK 0.414 0.715 0.618 0.582
QBench 0.513 0.926 0.246 0.562
MMBench 0.465 0.920 0.252 0.546
Gold (Ours) 0.579 0.833 0.557 0.656
@@
Dimensional Diversity. Dimensional diversity ($D_{\text{div}}$) measures how distinct and non-overlapping the latent ability dimensions are. It is computed as the inverse of the highest Heterotrait–Monotrait (HTMT) ratio across all construct pairs. Benchmarks like SEEDBenchV2 (0.605), MMEBench (0.549), and our Gold (0.579) exhibit strong discriminant validity across dimensions, reflecting diverse and well-separated cognitive constructs. In contrast, MVBench (0.094) shows the lowest $D_{\text{div}}$, suggesting highly entangled constructs that may compromise diagnostic clarity.
Task Contribution. Task contribution ($TC$) assesses the average magnitude of standardized loadings for all indicators across their respective constructs. High $TC$ values indicate that tasks meaningfully shape latent abilities with low redundancy. Most benchmarks perform well on this metric, including QBench (0.926), MMBench (0.920), and MMEBench (0.861). Our Gold benchmark (0.833) maintains a high $TC$ while avoiding excessive task overlap. In contrast, MVBench again underperforms (0.584), indicating weaker task-level informativeness and possible construct inflation due to redundant indicators.
Indicator Validity. Indicator validity ($D_{\text{valid}}$) evaluates how independent task indicators are in terms of multicollinearity. Computed as the inverse geometric mean of their VIFs, higher scores imply stronger statistical uniqueness across tasks. BLINK (0.618), Gold (0.557), and MVBench (0.522) exhibit the highest indicator validity, suggesting that their tasks are statistically independent and less redundant. Meanwhile, SEEDBenchV2 (0.081) and QBench (0.246) show weak indicator validity, indicating significant multicollinearity among tasks that may hinder interpretability.
Correlation Analysis We evaluated the Pearson correlation between benchmark scores and human evaluation results. Among all compared benchmarks, Gold (ours) achieves the highest correlation (0.7359), demonstrating its superior alignment with human judgment. MMMU also shows strong consistency (0.702), whereas BLINK and MVBench exhibit weaker correlations (0.2746 and 0.369), suggesting that certain benchmarks may not effectively reflect real-world reasoning capabilities.
5 Conclusion
@@
Appendix A Technical Appendices and Supplementary Material
A.1 Related Work
A.1.1 MLLM Benchmarks
A variety of benchmark datasets have been proposed to evaluate the capabilities of multimodal large language models (MLLMs), covering visual perception, language understanding, external knowledge integration, and complex reasoning. Early datasets such as COCO [37], VQA [38], and Visual7W [39] mainly assess perception and basic question answering abilities. Subsequent benchmarks like GQA [40] and VCR [41] emphasize compositional and commonsense reasoning, while OKVQA [42] and TextVQA [43] introduce challenges involving external knowledge and OCR-based understanding. ScienceQA [44] further targets scientific reasoning in educational contexts. As MLLMs develop, these earlier benchmarks are no longer sufficient to capture their increasingly diverse capabilities. Recent efforts, such as MMBench [25], OmniBench [45], and UniBench [46], aim to provide broader evaluations that include mathematics, medical diagnosis, chart understanding, and instruction following. However, these benchmarks typically rely on heuristic or empirical task categorization, leading to overlapping definitions, inconsistent classification, and limited diagnostic granularity, thus hindering efficiency analysis of model capabilities.
A.1.2 Evaluation for Benchmarks
A.2 SEM-based Evaluation Framework Supported Paradigms
SEM can be categorized into formative models and reflective models based on the relationship between observed indicators and latent variables. In a reflective model, the latent variable is considered the source of variation in the observed variables, which are reflections of the latent variable. In contrast, a formative model assumes that the latent variable is "formed" by multiple observed indicators, and its value depends on the composition of these indicators.
In the evaluation framework for MLLMs, we focus on the model’s performance across various specific tasks, and these task scores collectively determine our judgment of a particular latent ability. In other words, a certain ability does not preexist and dominate the performance of individual tasks, but is instead constituted by the aggregation of task performance. This logic of "multiple observed indicators forming a latent variable" aligns clearly with the definition of a formative measurement model.
Therefore, in constructing the SEM for evaluating MLLM capabilities, it is more appropriate to employ a formative measurement model. Unlike reflective models that assume latent variables give rise to observed indicators, the formative approach better captures the nature of MLLM evaluation, wherein diverse task-specific scores contribute to the formation of underlying cognitive abilities. As shown in Figure 1, the formative measurement model specifies that each latent ability dimension $\xi_{j}$ is formed as a linear combination of its automatic task-level indicators, i.e., $\xi_{j}=\sum_{m_{j}}\lambda_{x_{m_{j},j}}x_{m_{j},j}$, where $x_{m_{j},j}$ denotes the performance score of the MLLM on the $m_{j}$-th task corresponding to the $j$-th ability, and $\lambda_{x_{m_{j},j}}$ represents the factor loading that quantifies the contribution of each task to the latent construct. Furthermore, the overall capability of the MLLM, denoted by $Y$, is conceptualized as a second-order latent construct synthesized from the individual ability dimensions, expressed as $Y=\sum_{i}\omega_{i}\xi_{i}$, where $\omega_{i}$ denotes the weight assigned to the $i$-th ability dimension. This hierarchical formulation enables a comprehensive and theoretically grounded assessment of MLLM performance, accounting for both granular ability-level indicators and the integrated overall capability.
A.3 More details about Hierarchical Structure Based on Piaget’s Developmental Theory
Concretely, inspired by the major stages of Piaget’s theory, we define three core layers of MLLM capabilities:
  • Perception Layer (Sensorimotor Stage): This layer captures low-level perceptual and sensory processing abilities, including visual parsing, object recognition, attribute identification, and vision-language spatial alignment. It tests whether the model can robustly process and ground raw multimodal inputs.
  • Memory Layer (Preoperational Stage): This layer focuses on symbolic representation and contextual memory, encompassing tasks such as long-range dependency tracking, metaphor comprehension, and situational recall. It assesses the model’s ability to retain, represent, and retrieve information in dynamic or abstract contexts.
  • Reasoning Layer (Concrete and Formal Operational Stages): This layer evaluates high-level cognitive capabilities, such as causal inference, multi-hop reasoning, mathematical problem-solving, and abstract cross-modal generalization. Tasks in this category are designed to test whether the model can perform complex, multi-step reasoning and apply knowledge flexibly to novel problems.
A.4 Limitations

### v2

@@ V-B Main Results
In this section, we report the evaluation outcomes of various open-source and closed-source MLLMs on the Gold benchmark under the Main Experiment setting. Table IX presents the detailed results, where models are assessed across three cognitive layers—Perception, Memory, and Reasoning—further subdivided into interpretable subtasks such as color recognition, counting, OCR, basic medical knowledge (BMK), and domain-specific logical reasoning. BMK and CS denote Basic Medical Knowledge and Computer Science, respectively. Overall scores are computed by weighting capability-level scores according to task contributions.
Overall Performance. Among closed-source models, GPT-4o achieves the highest overall score of 216.3, excelling particularly in Perception (Color 92.5, Count 92.5, OCR 96.3) and Memory tasks (Artwork 72.6, Landmark 91.0, BMK 73.3). Gemini-2.0 variants (Pro: 212.9, Flash: 215.5) also perform strongly, with Gemini-2.0-Flash achieving the highest OCR score (96.3) among closed-source models, while Gemini-2.0-Pro leads in Memory (Landmark 92.8, BMK 83.3). Among open-source models, InternVL3-78B reaches 221.1 overall, surpassing all other models including GPT-4o, while Qwen2.5-VL-72B achieves 207.2 overall, showing substantial improvements over smaller variants (3B: 160.1, 7B: 195.9).
Capability-Level Comparison. Perception tasks are consistently the strongest across all models. For example, InternVL3-78B scores 91.7 (Color), 85.0 (Count), and 96.3 (OCR). DeepSeek models, though smaller (1B–27B), also show solid Perception performance, with DeepSeek-VL2 achieving 90.0 (Color) and 86.1 (Landmark), demonstrating that even lightweight open-source models can handle low-level visual understanding. Memory tasks show more variability: InternVL3-78B achieves 92.9 on Landmark but only 83.3 on BMK, while Qwen2.5-VL-72B scores 79.8 on Landmark and 73.3 on BMK. DeepSeek-VL2-small (16B) reaches 84.9 on Artwork and 53.3 on BMK, highlighting a gap between small-to-medium and large models. Reasoning tasks remain the most challenging across all families, with GPT-4o scoring 70.0 in Math and 80.0 in Economics, InternVL3-78B scoring 66.7 and 80.0 respectively, and DeepSeek models struggling further, e.g., DeepSeek 1B scoring only 16.7 in Math.
Model Size and Performance. Performance generally scales with model size. InternVL3 demonstrates steady gains from 161.8 (2B) to 221.1 (78B). Qwen2.5-VL improves from 160.1 (3B) to 207.2 (72B), while DeepSeek scales from 118.2 (1B) to 173.8 (27B), showing that even smaller architectures benefit from parameter growth. However, Reasoning subfields such as Electronics, Math, and Economics remain bottlenecks for all families.
V-C SEM-based Benchmark Evaluation Result Analysis
@@
Figure 8: Diagnostic metrics for existing MLLM benchmarks across four dimensions.
As shown in Fig. 8, Gold achieves the best balance among the three evaluation dimensions, with the highest overall score (2.625). It combines broad cognitive coverage, strong task informativeness, and robust statistical validity, demonstrating the strength of theory-driven benchmark restructuring under the SEM-based framework. In contrast, some existing benchmarks show localized strengths but lack holistic consistency. Moreover, Gold shows the highest alignment with human ratings (Pearson $r=0.7359$), reflecting its superior interpretability and evaluation reliability.
Dimensional Diversity. $D_{\mathrm{div}}$ measures distinct and non-overlapping the latent ability dimensions are. It is computed as the inverse of the highest Heterotrait–Monotrait (HTMT) ratio across all construct pairs. Benchmarks like SEEDBenchV2 (0.605), MMEBench (0.549), and our Gold (0.579) exhibit strong discriminant validity across dimensions, reflecting diverse and well-separated cognitive constructs. In contrast, MVBench (0.094) shows the lowest $D_{\mathrm{div}}$, suggesting highly entangled constructs that may compromise diagnostic clarity.
Task Contribution. $\mathrm{TC}$ assesses the average magnitude of standardized loadings for all indicators across their respective constructs. Higher $\mathrm{TC}$ values suggest that tasks substantially shape latent abilities while minimizing redundancy. Most benchmarks perform well on this metric, including QBench (0.926), MMBench (0.920), and MMEBench (0.861). Our Gold benchmark (0.833) maintains a high $\mathrm{TC}$ while avoiding excessive task overlap. In contrast, MVBench again underperforms (0.584), indicating weaker task-level informativeness and possible construct inflation due to redundant indicators.
Indicator Validity. $D_{\mathrm{valid}}$ evaluates how independent task indicators are in terms of multicollinearity. Computed as the inverse geometric mean of their VIFs, higher scores imply stronger statistical uniqueness across tasks. BLINK (0.618), Gold (0.557), and MVBench (0.522) exhibit the highest indicator validity, suggesting that their tasks are statistically independent and less redundant. Meanwhile, SEEDBenchV2 (0.081) and QBench (0.246) show weak indicator validity, indicating significant multicollinearity among tasks that may hinder interpretability.
@@
Correlation Analysis We evaluated the Pearson correlation between benchmark scores and human evaluation results, as shown in Fig. 9. Among all compared benchmarks, Gold (ours) achieves the highest correlation (0.7359), demonstrating its superior alignment with human judgment. MMMU also shows strong consistency (0.702), whereas BLINK and MVBench exhibit weaker correlations (0.2746 and 0.369), suggesting that benchmarks may not effectively reflect real-world reasoning capabilities.
VI Conclusion and Future Work